OpenStudy (anonymous):

how do i find the slope of this complex number 2401i?

OpenStudy (amistre64):

id say by plotting it, but i dont recall ever hearing about the slope of a complex number before

OpenStudy (amistre64):

if anything its the tangent of b/a

OpenStudy (anonymous):

slope??

OpenStudy (anonymous):

would taht result in a 0 slope?

OpenStudy (amistre64):

no, the slope of a vertical line is simply undefined

OpenStudy (anonymous):

not slope i mean degrees

OpenStudy (anonymous):

you are on the imaginary axis. there is no slope, and in any case a complex number is a point in the complex plain, not a line, so it has not slope

OpenStudy (amistre64):

your on the postive i axis if anything ....

OpenStudy (anonymous):



OpenStudy (amistre64):

like that

OpenStudy (anonymous):

so how would i apply the DeMoivre's theorem is i need the cube root of this imaginary number?

OpenStudy (amistre64):

wouldnt you just fill in the parts that it requires?

OpenStudy (anonymous):

2401i= 0+2401i there fore according to DeMoivre's theorem 2401^1/4(cos0+i sin0)

OpenStudy (anonymous):

but that says its not correct

OpenStudy (amistre64):

i think the 0 is in the wrong spot

OpenStudy (anonymous):

hmm you mean as a part of the standard form of the imaginary number 0+2401i?

OpenStudy (amistre64):

z^n= r^n (cos nθ + i sin nθ) right?

OpenStudy (anonymous):

correct for the trigonometric form of a complex number

OpenStudy (amistre64):

\[z^{1/3}=2401^{1/3}(cos(\frac{pi}{6})+isin(\frac{pi}{6}))\] if im reading this right

OpenStudy (anonymous):

thanks for your patience. the question reads what are the fourth roots of 2401i.

OpenStudy (amistre64):

oh, cause i read "so how would i apply the DeMoivre's theorem is i need the cube root of this imaginary number?" :)

OpenStudy (anonymous):

oh who knew? \[7^4=2401\]!!

OpenStudy (anonymous):

so I need the 4 root of the complex number which needs to be in trig form but i cannot figure out theta

OpenStudy (amistre64):

theta = pi/2

OpenStudy (anonymous):

why is it pi/2?

OpenStudy (anonymous):

@amistre i wolframed it. had no idea that it is \(7^4\)

OpenStudy (amistre64):

plot 2401i in the complex plane; youll see that its sitting right there n the "y" axis

OpenStudy (amistre64):

lol, that one of the most common 4ths too aint it?

OpenStudy (anonymous):

ok so x is zero and it only goes up the imaginary axis hence pi/2

OpenStudy (anonymous):

guess that was the clue that it was 4th roots they were after

OpenStudy (amistre64):

i havent even taken my finals yet and ive got 5 As regardless ..... the 6th one is prolly an A, if not its a B+

OpenStudy (anonymous):



OpenStudy (anonymous):

any dancing this semester?

OpenStudy (anonymous):

yes! that is the graph, how did you get those angles?

OpenStudy (amistre64):

not this semester ...

OpenStudy (anonymous):

divided \(\frac{\pi}{2}\) in by 4, got \(\frac{\pi}{8}\) and then divided the cirlce up into 4 equal parts taking all math or something more fun?

OpenStudy (amistre64):

3 maths, religion, science, and computers

OpenStudy (anonymous):

one of these things is not like the others....

OpenStudy (anonymous):

ok thanks a lot for your help guys

OpenStudy (amistre64):

yw

OpenStudy (anonymous):

\[7\left(\cos(\frac{\pi}{8})+i\sin(\frac{\pi}{8})\right)\] etc

OpenStudy (anonymous):

awesome thanks